# mpdframe frametitlebackgroundcolor =black!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, skipabove =nermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

### v1

@@ V-D Mitigating Miscompilations
For bug M1, Galápagos produces 10 equivalent variants in the same-language configuration. It then assembles an 11-Version implementation of the function which triggers the miscompilation fault. The key result here is in the ability of the 11-Version function to let developers know that something wrong has happened. The original function compiles, and when it executes, produces a value. However, the value is wrong, because of the miscompilation bug, but there is no warning or check to let the developers know about this wrong value. On the other hand, the 11-Version function compiles, and when it executes it crashes because the results returned by each version are not consistent. In this case, the program does not silently introduce a wrong value and the developer can act upon this. Consider the control flow graphs in Figure 11. The graph on the left represents the execution of a single-version implementation of the M1 function. The miscompiled code could result in the program silently choosing the wrong execution path, with potentially dangerous results. The graph on the right represents the execution of an N-Version implementation of the M1 function. The miscompiled code executes and returns silently, however, if at least any of the other 10 versions returns a distinct value, the program will exit, thus avoiding wrong execution and informing the program operator that something wrong has occurred. This behavior is similar for the N-Version implementations that Galápagos produced for M2 and M3, with N values of 11 and 3, respectively.
Upon manual inspection of the execution of the resulting binaries, we find that none of the equivalent variants are affected by the miscompilation bugs. The crashes are triggered by the inconsistent return values of the original function, and the return function of one of the corresponding variants.
@@ VI Discussion
In this section, we discuss the key design decisions that fix the scope of Galápagos, as well as the threats to the validity of our experiments.
VI-A Threats to Validity
@@ VII Related Work
Automated Variant Synthesis: Singh et al. [38] propose the automatic creation of language-diverse, formally verified program variants from a high-level specification. They focus on fault-tolerance of distributed systems by executing each variant following the replicated state machine paradigm. However, no prototype implementation, experimental methodology, or results are provided. Pelofske et al. [18] present a thorough analysis of function variant generation with LLMs. This study is focused on four SHA-1 functions, and provides a deep insight into the diversity and correctness of LLM-generated variant functions. Xu et al. [17] describe a variant generation mechanism based on LLVM-IR obfuscation. Their goal is to create explicitly non-equivalent variants, which are resistant to replication of known tampering mechanisms. While these works aim at automating variant creation, and to different extents, correctness verification, the key novelty of our work is the focus on hardening of binaries via N-Version function execution.
@@
LLMs code correctness: Yang et al. [22] present a C-to-Rust verified translation tool. It works by using static methods to construct a known equivalent translation, and LLMs to create more idiomatic translations. The latter are then checked for equivalence against the former using an off-the-shelf Rust tool. In the same vein, Bathia et al. [39] describe a tool to produce verified code transpilations. This is achieved by leveraging intermediate representations and formal equivalence checkers. While the mentioned works address the issue of formally proving LLM-generated code equivalence, their scope is not concerned with software reliability. Du et al. [40] present a benchmark for solutions to programming tasks via LLMs. This benchmark is designed to measure how efficient are the generated solutions in terms of algorithmic complexity. The experiments show that LLMs are able to produce code that solves the tasks while having diverse efficiency scores. Zhang et al. [28] introduce a benchmark to quantify the correctness of code produced by an LLM. The experiments evidence that LLMs have the capability to correctly solve programming tasks to a large extent. However, while the data provided by these works gives a detailed overview of the correctness and efficiency of code produced by LLMs, the evaluations are test-based, and not formally verified against a specification.
@@
Code Transplantation and Recombination: As part of Galápagos’ Harnessing pass, IR code from newly generated programs is inserted into the project that is being hardened. This process is similar to the code transplantation mechanisms described in the literature. This process has been studied in the context of automated program repair [41], feature transfer between systems [42, 43, 44, 45], and code clone analysis [46]. The Harnessing pass is also similar to recombination approaches, where different versions of systems are created by manipulating sections of readily available codebases, at lower-than-source levels [47, 48]. While closely related, these works are not concerned with the challenges of the two first passes of Galápagos, namely, automatic diversification and correctness checking.
@@
Addressing Miscompilations: The works of Eide et al. [26] and Le et al. [49] present approaches to discover miscompilation bugs via stochastic exploration of the program space. With the same goal, Tu et al. [50] describe a mechanism with guided LLM prompt production approach, which allows producing bug-triggering program mutations. While these approaches seek to actively discover miscompilation bugs, Galápagos takes a program hardening, as it does not target compilers directly, but protects critical application code against potential miscompilation.
@@ VIII Conclusion
In this paper, we have introduced the original idea of automated N-Version programming with LLMs. The key idea is to exploit the search power and coding capabilities of LLMs, within the scope of N-Version programming. We presented Galápagos, a system which allows for automatic N-Version programming, with formal guarantees. We evaluated its performance in three dimensions: variant correctness, code diversity, and ability to counteract miscompilation bugs.
Our experiments show that: (1) Galápagos is able to produce code variants with LLMs, that can be proven to be equivalent, even when these variants are written in a different programming language. (2) Galápagos is able to produce code variants which are diverse both on disk and at execution-time, and; (3) the Galápagos variants can be harnessed to harden critical sections of software against the dangerous class of miscompilation bugs. Our experiment clearly show that the cost of producing N-Version software can be reduced by the combined use off-the-shelf generative language models.
We identify two future angles for this novel research direction. Galápagos relies on N-of-N voting, meaning that all variants must produce the same output for the program to continue execution. However, previous works in the N-Version programming line have proposed alternative voting mechanisms [51]. Galápagos can be extended to support different voting mechanisms, such that voting can be tailored to the context and specific requirements of the application. For instance, a threshold voting (k-of-N) mechanism, would allow Galápagos to provide similar hardening enhancements, while also hiding undesired behavior, and continue with execution.

### v2

@@ V-D Mitigating Miscompilations
For bug M1, Galápagos produces 10 equivalent variants in the same-language configuration. It then assembles an 11-Version implementation of the function which triggers the miscompilation fault. The key result here is in the ability of the 11-Version function to let developers know that something wrong has happened. The original function compiles and, when it executes, produces a value. However, the value is wrong because of the miscompilation bug, but there is no warning or check to let the developers know about this wrong value. On the other hand, the 11-Version function compiles, and when it executes, it crashes because the results returned by each version are not consistent. In this case, the program does not silently introduce a wrong value and the developer can act upon this. Consider the control flow graphs in Figure 11. The graph on the left represents the execution of a single-version implementation of the M1 function. The miscompiled code could result in the program silently choosing the wrong execution path, with potentially dangerous results. The graph on the right represents the execution of an N-Version implementation of the M1 function. The miscompiled code executes and returns silently; however, if at least any of the other 10 versions returns a distinct value, the program will exit, thus avoiding wrong execution and informing the program operator that something wrong has occurred. This behavior is similar for the N-Version implementations that Galápagos produced for M2 and M3, with N values of 11 and 3, respectively.
Upon manual inspection of the execution of the resulting binaries, we find that none of the equivalent variants are affected by the miscompilation bugs, only the original function is. As a result, the program termination is triggered by the inconsistent return values between the original function and the return values of one of the corresponding variants.
@@
V-E Performance Overhead
Table VII shows a summary of the overhead introduced by the N-Version functions when compared to the reference functions. The column "Baseline Runtime" shows the measurement taken for each function as described in subsection IV-F. For both same-language and cross-language configurations, "N" represents the number of versions used in the maximally diverse N-version function; "N-Version runtime" shows the execution time for the N-version function; and finally "Normalized Overhead" shows the average overhead for a N-version function with one single-variant (i.e. a 2-version function).
The average normalized overhead is expected to be 2x for a 2-version function, which is what we observe on the last average row. The observed maximum and minimum overheads are 3.75x and 1.04x respectively. The minimum value can be considered an outlier, and it is introduced by a loop-unroll compiler optimization in the Go variant of a C function. We note that it is also possible to design parallel execution harnesses in time-critical contexts.
{mdframed}
[style=mpdframe, frametitle=Answer to RQ4] Our experiments show that the N-Version functions generated by Galápagos introduce the expected overhead. After executing the N-Version functions, we observe that the runtime is increased by 2x for a 2-version function, and grows linearly by adding more variants to the assembly. This experiment demonstrates that the harness overhead overhead is negligible.
Function Name Baseline runtime ($\mu$s) Same-language diversification (C) Cross-language diversification (C+Go)
N N-Version runtime ($\mu$s) Normalized Overhead N N-Version runtime ($\mu$s) Normalized Overhead
alaw_to_s16 4 606 2 17 256 3.75x 7 38 363 2.22x
iec958_parity 83 623 5 269 676 1.56x 5 95 393 1.04x
add 2 486 10 36 362 2.51x 2 5 879 2.36x
ulaw_to_s16 4 369 10 54 707 2.28x 11 52 961 2.11x
val_seg 4 163 3 12 544 2.01x
flac_get_max_frame_size 6 222 5 30 622 1.98x
mix 9 067 7 60 226 1.94x
weight 6 256 11 59 428 1.85x
icbrt64 64 412 8 700 098 2.41x
BitDeinterleave 1 745 4 7 309 2.06x
BitInterleave 1 848 2 4 293 2.32x
_booth_recode_w5 4 599 3 13 757 2.0x 3 11 070 1.7x
fpr_half 2 895 3 12 139 2.6x 2 5 833 2.01x
fpr_lt 5 615 2 13 796 2.46x
modp_montymul 5 543 3 20 754 2.37x
modp_norm 2 793 3 8 210 1.97x
int16_nonzero_mask 3 872 4 15 601 2.01x 9 25 391 1.69x
barrett_reduce 3 095 11 67 948 3.1x
ctz 212 299 10 866 981 1.34x 7 655 523 1.35x
int16_t_negative_mask 2 754 8 25 538 2.18x 6 16 344 1.99x
int16_t_nonzero_mask 4 060 8 27 467 1.82x 6 18 922 1.73x
montgomery_reduce 2 472 5 18 241 2.59x 9 17 304 1.75x
fBlaMka 3 356 10 38 936 2.18x
Avg. Overhead 2.18x 1.99x
TABLE VII: Runtime measurement of N-Version functions.
@@ VI Discussion
In this section, we discuss Galápagos in the larger context of software development, as well as the threats to the validity of our experiments.
VI-A Integration into Existing Software Development Practices
The integration of Galápagos into an existing software development project requires: (1) identifying critical sections of code with the highest reliability standards, and; (2) isolating them into pure functions. Developers would need to integrate Galápagos in the build pipeline for those functions. The pipeline would automatically generate variants, validate their correctness, and insert the N-Version functions into the final build. At deployment time, the N-version functions would terminate execution in case of inconsistencies detected by diversification. This is valid in IoT [38]. We are also aware that companies providing security critical software for the military have such build pipelines with other kinds of diversification.
VI-B Threats to Validity
@@ VII-A Automated Variant Synthesis
Singh et al. [39] propose the automatic creation of language-diverse, formally verified program variants from a high-level specification. They focus on fault-tolerance of distributed systems by executing each variant following the replicated state machine paradigm. However, no prototype implementation, experimental methodology, or results are provided. Pelofske et al. [18] present a thorough analysis of function variant generation with LLMs. This study is focused on four SHA-1 functions, and provides a deep insight into the diversity and correctness of LLM-generated variant functions. Xu et al. [17] describe a variant generation mechanism based on LLVM-IR obfuscation. Their goal is to create explicitly non-equivalent variants, which are resistant to replication of known tampering mechanisms. While these works aim at automating variant creation, and to different extents, correctness verification, the key novelty of our work is the focus on hardening of binaries via N-Version function execution.
VII-B LLMs code correctness
@@
Pan et al. [40] identify the correctness of LLM-generated code as an important challenge to address. In this vein, Kessel and Atkinson [41] present LASSO, a platform for differential execution of large sets of code versions and tests. Compared to this work Galápagos goes one step further and automates the creation of N-Version programs, in addition to generation and testing of code variants. Du et al. [42] present a benchmark for solutions to programming tasks via LLMs. This benchmark is designed to measure how efficient are the generated solutions in terms of algorithmic complexity. The experiments show that LLMs are able to produce code that solves the tasks while having diverse efficiency scores. Zhang et al. [28] introduce a benchmark to quantify the correctness of code produced by an LLM. The experiments evidence that LLMs have the capability to correctly solve programming tasks to a large extent. However, while the data provided by these works gives a detailed overview of the correctness and efficiency of code produced by LLMs, the evaluations are test-based, and not formally verified against a specification.
VII-C Code Generation and Code Translation
Code generation and translation with LLMs are very active research areas [43, 44]. Yang et al. [22] present a C-to-Rust verified translation tool. It works by using static methods to construct a known equivalent translation, and LLMs to create more idiomatic translations. The latter are then checked for equivalence against the former using an off-the-shelf Rust tool. Bathia et al. [45] describe a tool to produce verified code transpilations. This is achieved by leveraging intermediate representations and formal equivalence checkers. Yang et al. [46] present UniTrans, a framework to generate and test LLM code translation. UniTrans uses the original code to further generate test cases for the resulting translation, and leverage them to iteratively repair the translation in case of failure. While the mentioned works address the issue of validating correctness of LLM-generated code equivalence, their scope is not concerned with software reliability.
VII-D Code Transplantation and Recombination:
@@
As part of Galápagos’ Harnessing pass, IR code from newly generated programs is inserted into the project that is being hardened. This process is similar to the code transplantation mechanisms described in the literature. This process has been studied in the context of automated program repair [47], feature transfer between systems [48, 49, 50, 51], and code clone analysis [52]. The Harnessing pass is also similar to recombination approaches, where different versions of systems are created by manipulating sections of readily available codebases, at lower-than-source levels [53, 54]. While closely related, these works are not concerned with the challenges of the two first passes of Galápagos, namely, automatic diversification and correctness checking.
VII-E Addressing Miscompilations
@@
The works of Eide et al. [26] and Le et al. [55] present approaches to discover miscompilation bugs via stochastic exploration of the program space. With the same goal, Tu et al. [56] describe a mechanism with guided LLM prompt production approach, which allows producing bug-triggering program mutations. While these approaches seek to actively discover miscompilation bugs, Galápagos takes a program hardening, as it does not target compilers directly, but protects critical application code against potential miscompilation.
@@ VIII Conclusion
In this paper, we have introduced the original idea of automated N-Version programming with LLMs. The key idea is to exploit the search power and coding capabilities of LLMs, within the scope of N-Version programming. We presented Galápagos, a system that allows for automatic N-Version programming, with formal guarantees. We evaluated it in four dimensions: variant correctness, code diversity, the ability to counteract miscompilation bugs, and introduced overhead.
Our experiments show that: (1) Galápagos is able to produce code variants with LLMs, that can be proven to be equivalent, even when these variants are written in a different programming language. (2) Galápagos is able to produce code variants which are diverse both on disk and at execution-time; (3) the Galápagos variants can be harnessed to harden critical sections of software against the dangerous class of miscompilation bugs, and; (4) the N-Version functions generated by Galápagos introduce a performance overhead that stays within reasonable and expected limits. Finally, our experiments clearly show that the production of N-Version software can be automated by the combined use of off-the-shelf tools and generative language models.
We identify two future angles for this novel research direction. Galápagos relies on N-of-N voting, which means that all variants must produce the same output for the program to continue to execute. However, previous works in the N-Version programming line have proposed alternative voting mechanisms [57]. Galápagos can be extended to support different voting mechanisms, so that voting can be tailored to the context and specific requirements of the application. For instance, a threshold voting mechanism (k-of-N) would allow Galápagos to provide similar hardening enhancements, while also hiding unwanted behavior, and continue with execution.